One point is scored for each successful shot. So the player can end up with 0 points(by missing the first shot), q point (by making first shot but not second shot), or 2 points (by making both shots).
Terry is a basketball player who has shown over a period of time that she makes about 60% of the free throws she attemps.
Now suppose that Terry is in a one-on-one situation. How many points is she most likely to score: 1, 2, 3?

Answer by longjonsilver(2297)

P(scores) = 0.6, so that P(misses) = 0.4

P(score) and P(score again) = 0.6*0.6 = 0.36
P(score) and P(miss) = 0.6*0.4 = 0.24
P(miss) and P(score) = 0.4*0.6 = 0.24
P(miss) and P(miss again) = 0.4*0.4 = 0.16

so we can now think about the possible situations:
P(no score) = 0.16
P(score=1) = 0.24+0.24 = 0.48
P(score=2) = 0.36

So, most likely to score 1 point.

jon